OpenStudy (firejay5):

4.You push a friend sitting on a swing with a force of 210N. She accelerates at a rate of 4 m/s^2. A. What is the mass of your friend? B. What is your friend's weight?

OpenStudy (firejay5):

A. M = F/A M = 210N/4m/s^2 M = 52.5kg or 53kg B. F = M * A F = 53kg * 4m/s^2 F = 212N

OpenStudy (firejay5):

@caozeyuan @Noodles24

OpenStudy (caozeyuan):

B is wrong

OpenStudy (caozeyuan):

weight is gravitational force

OpenStudy (caozeyuan):

and G=mg, g=9.81ms^-2

OpenStudy (firejay5):

@caozeyuan What's the unit? 519.4 what???

OpenStudy (caozeyuan):

force in N, mass in kg, acceleration in ms^-2

OpenStudy (firejay5):

519.4 Newtons